OpenStudy (anonymous):

I really need help with these questions: what is the solution to (4 * 10^3) * (3 * 10^4) written in scientific notation? Jonny spent 1/8 of his allowance on video games. What is the fraction of his allowance expressed as a decimal? Find the product. 1/6 * 2/3

Vocaloid (vocaloid):

(4 * 10^3) * (3 * 10^4) first step is to add the exponents together so 10^3 times 10^4 = ?

OpenStudy (anonymous):

7?

OpenStudy (anonymous):

i guess

Vocaloid (vocaloid):

good so 10^3 times 10^4 = 10^7 now we look at the non-exponent parts 3*4 = ?

OpenStudy (anonymous):

12?

Vocaloid (vocaloid):

good so (4 * 10^3) * (3 * 10^4) = 12 * 10^7 last step is to rewrite in scientific notation, any ideas how to do that?

OpenStudy (anonymous):

no, not really

Vocaloid (vocaloid):

start by dividing 12 by 10 to get 1.2 then multiply 10^7 by 10 to balance things out what is 10^7 times 10?

OpenStudy (anonymous):

oh wait! I do know!

Vocaloid (vocaloid):

go ahead

OpenStudy (anonymous):

1200000.! is that correct? what do you think?

Vocaloid (vocaloid):

not scientific notation

Vocaloid (vocaloid):

scientific notation looks like this: X.X * 10^X where the X's are numbers

OpenStudy (anonymous):

oh woops the way I was taught is by using zeros

Vocaloid (vocaloid):

nah don't apologize let's go back to what i was saying before what is 10^7 times 10?

OpenStudy (anonymous):

100000000! is this correct?

OpenStudy (anonymous):

or...

Vocaloid (vocaloid):

in scientific notation please

Vocaloid (vocaloid):

10^7 * 10 = 10^____ fill in the blank

Vocaloid (vocaloid):

hint: add the exponents

OpenStudy (anonymous):

10^7 * 10 = 1.2^8

Vocaloid (vocaloid):

almost

Vocaloid (vocaloid):

1.2*10^8

Vocaloid (vocaloid):

ready for the next?

OpenStudy (anonymous):

yes

Vocaloid (vocaloid):

Jonny spent 1/8 of his allowance on video games. What is the fraction of his allowance expressed as a decimal? well, let's start off with an easier question what is 1/2 as a decimal?

OpenStudy (anonymous):

0.5

Vocaloid (vocaloid):

good! now, 1/4 is half of 1/2 so if 1/2 as a decimal is 0.5, what is 1/4 as a decimal?

OpenStudy (anonymous):

0.125

Vocaloid (vocaloid):

well uh 1/4 as a decimal is 0.25 1/8 is 0.125

Vocaloid (vocaloid):

ready for the next?

OpenStudy (anonymous):

oh wait! I just noticed that the fraction I was supposed to put into that question was 2/3! oops!

Vocaloid (vocaloid):

oh that's fine

Vocaloid (vocaloid):

well, 1/3 is 0.3333(repeated) so what do you think 2/3 is?

OpenStudy (anonymous):

0.- 6 (the line there is supposed to be on top of 6)

Vocaloid (vocaloid):

good!

Vocaloid (vocaloid):

ready for the last?

OpenStudy (anonymous):

YES!!! XD

OpenStudy (anonymous):

I've never really been good with fractions though...

Vocaloid (vocaloid):

nah it's all good 1/6 * 2/3 start by multiplying the numerators (top numbers) 1*2 = 2 then multiply the denominators (bottom numbers) 6*3 = 18 then put 'em together to get 2/18 know how to simplify?

OpenStudy (anonymous):

well..... not rly

Vocaloid (vocaloid):

just divide the numerator and denominator by 2 2/2 = 1 18/2 = 9 so 2/18 becomes 1/9 and that's it

OpenStudy (anonymous):

WOW!!! Thanks a lot Vocaloid! You're so smart!!

Vocaloid (vocaloid):

np

OpenStudy (anonymous):

:D